1. A stock has a beta of 1.85, the expected return on the market is 8 percent, and the risk-free rate is 3.2 percent. What must the expected return on this stock be?

2. A portfolio has 85 shares of Stock A that sell for $36 per share and 100 shares of Stock B that sell for $28 per share
a. what is the portfolio weight of stock A?
b. what is the portfolio weight of stock B?

3.
Answer #1

As per HOMEWORKLIB RULES when there are more than one question then we have to answer first question.

Calculation of expected return:

Expected return= risk free rate+beta*(market return- risk free rate)

Expected return= 3.2+1.85*(8-3.2)

Expected return= 3.2+ 1.85*4.8= 3.2+8.88= 12.08

Expected return is 12.08 percent